OpenStudy (anonymous):

Match each percent equation with the accurate description of the equation. Match Term Definition 90= p% x 50 A) 50 is the whole 60 = 0.5 x b B) 50 is the part a = 0.9 x 50 C) 90percent of the whole 50 = 60% x b D) 50 percent of the whole

OpenStudy (anonymous):

C

OpenStudy (anonymous):

90=p%*50 --- A 60=0.5*b --- D a=0.9*50 --- C 50=60%*b --- B